# A Comprehensive Survey of Deep Transfer Learning for Anomaly Detection in Industrial Time Series: Methods, Applications, and Directions

### v1

@@ 5.2 Directions for anomaly detection solution design
How data preprocessing should be conducted is an open question. For industrial applications, some researchers consider directly using time series data as input for training to be inefficient, thus suggesting deriving or selecting features from time series data by statistical methods or human experience to decrease the complexity of the dataset dramatically. On the other hand, this crops a lot of potentially useful information, e.g., the time series trend. Some researchers use machine parameters as features of the manufacturing process instead of using process data collected by sensors (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022). Others try different transformations of raw time series data, a common way being to transform 1D time series data to 2D image data (Liao et al., 2021; Wang et al., 2021; Zabin et al., 2023) or transforming time domain signals otherwise into the frequency domain (Lukic et al., 2016). However, as large-scale computation power and storage become cheaper and more accessible, it is becoming increasingly common to use deep learning techniques to directly process time series data (Xu et al., 2019; Maschler et al., 2021).
Data augmentation
@@
Data augmentation is useful for deep learning models because it can help to prevent overfitting. For deep transfer learning, when a model becomes too closely adapted to the specifics of the source domain, it may not be able to generalize well to some examples in the task domain. One important technique is to acquire effective synthetic data, e.g., using a simulation process or model to explore potential anomalous conditions by simulating industrial processes under parameters that cannot yet be experienced in the real-world. High fidelity and reliability of simulation data can provide training data at low cost and mitigate the problem of insufficient samples for deep transfer learning (Xu et al., 2019). Another way to generate effective synthetic data is to use GANs. GANs are trained on normal data only to generate indistinguishable samples from which abnormal samples are distinguished during the testing stage of the overall anomaly detection system based on their deviating data distribution (Jiang et al., 2019). To increase the number of anomalous samples and thus the robustness of the anomaly detection model, the technique of adversarial perturbation known from computer vision (Goodfellow et al., 2015) can be used.
Data imbalance
@@
DNNs perform well when they are trained on balanced datasets. However, in practice, it is difficult to get sufficient anomalous data for anomaly detection tasks. For example, manufacturing process is usually in a healthy state due to the pre-designed and optimized operation. Several ways exist to address the imbalanced dataset for time series anomaly detection. To deal with data imbalance, one way is to oversample the minority class, i.e., to randomly replicate samples from the minority class to equalize the number of samples from each class in each batch. Synthetic Minority Over-sampling Technique is an advanced method that creates synthetic samples to force the decision region of the minority class to become more general (Chawla et al., 2002). This technique is widely used in anomaly detection tasks in industry (Ijaz et al., 2018; Mokhtari et al., 2021). Apart from oversampling, the resampling strategy is also frequently used to assign a higher probability to abnormal samples and evenly select the same amount of samples from both classes in each batch. Finally, a weighted loss can be used that balances the loss for the abnormal and normal class in supervised anomaly detection (Buda et al., 2018).
@@ 5.3 Directions for deep transfer learning implementation
(1) Limited data availability: If the amount of data available for a specific task is limited, pre-training on related source data can learn general features that can be transferred to the specific learning task in the target domain. (2) Similar domains: deep transfer learning is well suited for similar source and target domains. (3) Limited resources (time and compute): Using parameter transfer here is recommended if pre-trained models exist.
@@
(1) Different tasks: If the target learning task is vastly different from the source learning task, deep transfer learning may not be appropriate. For example, if one wants to train a model for natural language processing on a new dataset, using a pre-trained model that has been trained on image data will not be useful. (2) High domain shift: If there is a large difference between the source and the target domain, deep transfer learning may not be effective. This can happen when the data distributions, features, or labels are vastly different. (3) Abundance of labeled data: If there are enough data for the new task, it may be more effective to train a model from scratch (Liang et al., 2018).
@@
We suggest selecting the model architecture mainly depending on the data size and label availability, starting from a relatively small network and moving gradually to more complex DNNs. It is important to effectively capture the temporal dependencies and extract respective features of time series data. LSTMs are used heavily for detecting temporal dependencies in time series data (Zabin et al., 2023; Panjapornpon et al., 2023; Xiong et al., 2018). CNNs are also effective in extracting time series features (Yao et al., 2022; Zgraggen et al., 2021). For semi-supervised settings, CNN-based auto-encoders are trained to reconstruct the original data (Serradilla et al., 2021).
Another aspect of deep transfer learning implementations is often the limited computing power of hardware platforms, such as embedded systems in industrial applications. Sensor data are typically acquired using resource-constrained edge processing devices that struggle with computationally intensive tasks, especially when updating a DNN model. One possible solution is federated learning, perhaps the most popular framework, mainly due to its feature of leveraging data while still preserving their privacy (Wang et al., 2018; Amiri and Gündüz, 2019). The technology enables a more collaborative approach to ML while preserving user privacy by storing data decentralized on distributed devices rather than on a central server. Combining deep transfer learning with federated learning is a promising and powerful combination in the abovementioned industrial applications.
@@
Foundation models such as transformers (Vaswani et al., 2017), diffusion models (Rombach et al., 2022) and SAM (Kirillov et al., 2023) demonstrate emerging properties such as in-context learning (Brown et al., 2020) and complex cross-modality conditioning. This is achieved by training complex and often auto-regressive models with massive amounts of data, although the precise mechanisms that lead to this are not well understood. Some of those models generalize to new settings and tasks, without an explicit element of transfer learning. Thus, the application of foundational models in industrial time series analysis has the potential to reduce and eventually eliminate the need to explicitly account for changes in the domain in modeling, and instead, the foundational models will provide the transfer capability. To not only detect anomalies but also identify failure modes and elicit an appropriate intervention, AI systems must have some form of understanding, a world model, or, in other words, the AI has to implicitly or explicitly model causal relations. Counterfactual inference incorporates causal relations between observations and interventions which allows making predictions of outcomes that were never seen during training (Vlontzos et al., 2023).
@@ 6 Conclusions
In this survey, we presented a comprehensive overview of (deep) transfer learning by defining transfer learning problem settings and categorizing the state-of-the-art deep transfer learning approaches. Equipped with this foundation, we selected representative examples of the landscape of fielded applications to provide practitioners with a guide to the field and possibilities of industrial time series anomaly detection. After carefully discussing open challenges, we gave practical directions for time series anomaly detection solution design and deep transfer learning implementation. We found that current applications focus on simple cases with simple datasets, neural network structures, and deep transfer learning schemes. Despite this, the survey suggests that deep transfer learning approaches have huge potential and promise for solving more complex and dynamic anomaly detection tasks in the industry. As the field is still in an early stage, more R&D is expected to fully realize the potential of deep transfer learning in increasingly complex settings. Future work should focus on developing robust transfer learning schemas and methods that can handle more complex and dynamic tasks. The following directions hold the greatest potential for future work:
Automatic selection of transferable features
(Long et al., 2015) It refers to methods for selecting and transferring only the relevant knowledge from the base model. This could involve the use of techniques such as selective fine-tuning and distillation to identify the most important features and knowledge learned from source domains (Yosinski et al., 2014; Ge and Yu, 2017).
Investing into more complex deep transfer learning schemes and DNN architectures
Most deep transfer learning approaches applied in industry focus on the parameter transfer approach as it is conceptionally the simplest and readily applicable by interdisciplinary teams without ML research experience. It seems promising to invest in testing more appropriate deep transfer learning approaches according to different use cases, such as mapping transfer, adversarial transfer, etc. The same applies to testing diverse DNN architectures besides straightforward ones.
@@
The data-centric approach focuses on improving ML models by ensuring high-quality labeled data (Stadelmann et al., 2022) using techniques such as re-labeling, re-weighting, or data augmentation (Luley et al., 2023). Currently, a human-in-the-loop solution is still needed, frameworks have been proposed to assist annotators with graph-based algorithms such as nearest neighbor graphs (Bai et al., 2021), decision trees (Liu et al., 2021), or factor graphs (Kang et al., 2022). Although these methods have proven to be effective, a more automated process is a goal for future research.
@@
To build robust AI approaches to solve time series anomaly detection in industry, only focusing on transfer learning will not be sufficient. Combinations with other ML approaches are needed in the future, such as continuous learning, meta-learning, and federated learning.
Acknowledgments
We would like to acknowledge Claudio Riginio for his constructive comments on an earlier draft of this paper and for his assistance with the illustrations. This work has been supported by Innosuisse grant 62174.1 IP-ENG “DISTRAL”.

### v2

@@ 5.3 Directions for anomaly detection solution design
How data preprocessing should be conducted is an open question. For industrial applications, some researchers contend that using raw time series data directly as input for training may not be the most efficient. Hence, they propose deriving or selecting features from time series data by statistical methods or human experience. This can significantly decrease the complexity of the dataset. On the other hand, this crops a lot of potentially useful information, e.g., the time series trend. To reduce the dimensionality, some researchers use machine parameters as features in the manufacturing process instead of the processing data collected by sensors [85, 10, 113, 114]. Others try different transformations of raw time series data, a common way being to transform 1D time series data to 2D image data [9, 15, 112] or transforming time domain signals otherwise into the frequency domain [39]. However, as large-scale computation power and storage become cheaper and more accessible, it is becoming increasingly common to use deep learning techniques to process time series data directly [13, 11].
Data augmentation by generative AI
@@
Data augmentation is useful for deep learning models because it can help to prevent overfitting. For deep transfer learning, when a model becomes too closely adapted to the specifics of the source domain, it may not be able to generalize well to some examples in the task domain. One important technique is to acquire effective synthetic data, e.g., using a simulation process or model to explore potential anomalous conditions by simulating industrial processes under parameters that cannot yet be experienced in the real world. High fidelity and reliable simulation data can provide training data at low cost and mitigate the problem of insufficient samples for deep transfer learning [13]. Another way to generate effective synthetic data is to use generative models, such as GANs. GANs are only trained on normal data to generate indistinguishable normal samples so that abnormal samples can be distinguished during the testing stage of the overarching anomaly detection system, as they deviate from the normal data distribution [134]. To increase the number of anomalous samples and thus the robustness of the anomaly detection model, the technique of adversarial perturbation known from computer vision [135] can be used.
Dealing with data imbalance
@@
DNNs perform well when they are trained on balanced datasets. However, in practice, it is difficult to get sufficient anomalous data for anomaly detection tasks. For example, the manufacturing process is usually in a healthy state due to the pre-designed and optimized operation. Several ways exist to address the imbalanced dataset for time series anomaly detection. One way is to oversample the minority class, e.g., by randomly replicating samples from the minority class to equalize the number of samples from each class in each batch. The Synthetic Minority Over-sampling Technique is an advanced method that creates synthetic samples to force the decision region of the minority class to become more general [136]. This technique is widely used in anomaly detection tasks in industry [137, 138]. Apart from oversampling, resampling strategies are frequently used to assign a higher probability to abnormal samples and evenly select the same amount of samples from both classes in each batch. Moreover, a weighted loss can be implemented to balance the loss between the abnormal and normal class in supervised anomaly detection [131].
@@ 5.4 Directions for deep transfer learning implementation
[…]
@@
[…]
@@
We recommend choosing the model architecture mainly based on data size and label availability, starting from a relatively small network and moving gradually to more complex DNNs. CNNs also effectively extract time series features [19, 111]. For semi-supervised settings, CNN-based auto-encoders are trained to reconstruct the original data [110]. It is important to effectively capture the temporal dependencies and extract features of time series data. LSTMs are extensively employed for this purpose, as they excel in detecting temporal dependencies in time series data [112, 21, 23].
[…]
@@
Foundation models like SAM [139] or others, using for example transformer architectures [48] or diffusion models [140], demonstrate emerging properties such as in-context learning [53] and complex cross-modality conditioning. This is achieved by training complex and often auto-regressive models with massive amounts of data, although the precise mechanisms that lead to this are not well understood. Some of those models generalize to new settings and tasks, without an explicit element of transfer learning. Thus, the application of foundation models in industrial time series analysis has the potential to reduce and eventually eliminate the need to explicitly account for changes in the domain within the modeling, by instead having the foundation model provide the transfer capability (see examples in Sec. 6-b). To not only detect anomalies but also identify failure modes, analyze root causes, and elicit an appropriate intervention, AI systems must implicitly or explicitly model causal relations. Counterfactual inference incorporates causal relations between observations and interventions, which allows predictions of outcomes never seen during training [141].
Another aspect of deep learning implementations is the limited computing power of hardware platforms, such as embedded systems in industry. Sensor data are typically acquired using resource-constrained edge processing devices that struggle with computationally intensive tasks, especially when training a DNN model. Federated learning stands out as a leading solution, with its ability to utilize data while preserving privacy [142, 143]. The technology enables a more collaborative approach to ML while preserving user privacy by storing data decentralized on distributed devices rather than on a central server. Combining deep transfer learning with federated learning is a promising and powerful combination in the abovementioned industrial applications.
@@ 6 Conclusions
In this survey, we presented a comprehensive overview of deep transfer learning by defining transfer learning problem settings and categorizing the state-of-the-art deep transfer learning approaches based on the surveyed papers. Then, we review and emphasize on investigating deep transfer learning approaches for time series anomaly detection in different industrial settings. Equipped with this foundation, we selected representative examples of the landscape of fielded applications to provide practitioners with a guide to the field and possibilities of industrial time series anomaly detection.
The main finding of this survey is that only a limited variety of deep transfer learning methods are employed in anomaly detection in industrial time series analysis – mainly simple ones. Almost all applications employ parameter transfer, arguably the most straightforward transfer approach. In its simplest implementation, it only involves fine-tuning a pre-trained model. Accordingly, the employed network architectures are simple, none of the reviewed research papers used advanced DNN building blocks like Transformer, which are common in computer vision and language modeling. We expect this type of architecture with suitable modifications and/or pre-trained parameters to spread to more niche fields. Despite this, the survey suggests that deep transfer learning approaches have huge potential and promise for solving more complex and dynamic anomaly detection tasks in industry. As the field is still in an early stage, more R&D is expected to fully realize the potential of deep transfer learning in increasingly complex settings.
In the end, we highlight the importance of considering feasibility, reliability, explainability, and real-time data stream when designing a transfer learning system for time series anomaly detection. After carefully discussing open challenges, we gave practical directions for time series anomaly detection solution design and deep transfer learning implementation. In our view, the following directions hold the greatest potential for future work:
Automatic selection of transferable features [57]
It refers to methods for selecting and transferring only the relevant knowledge for the new tasks from the base model. This could involve the use of techniques such as selective fine-tuning and distillation to identify the most important features learned from source domains [30, 144].
Investing into advanced deep transfer learning schemes and DNN models
[…]
@@
The data-centric approach focuses on improving ML models by ensuring high-quality labeled data [147] using techniques such as re-labeling, re-weighting, or data augmentation [148]. Currently, a human-in-the-loop solution is still needed. Frameworks have been proposed to assist annotators with graph-based algorithms such as nearest neighbor graphs [84], decision trees [149], or factor graphs [150]. Although these methods have proven to be effective, a more automated process is a goal for future research.
Leveraging generative AI
Generative models like GANs and diffusion models can generate synthetic time series data, making them valuable for data augmentation. Augmenting the original data with synthetic samples can enhance the deep learning models’ robustness, especially in real-world applications where target data are limited. These models can also be leveraged to examine anomalies and generate anomalies to help alleviate the imbalance within the data [151].
@@
To develop robust AI solutions for time series anomaly detection in the industry, relying solely on transfer learning is insufficient. Future strategies should integrate other ML approaches, including continuous learning, meta-learning, and federated learning.